# EMD: Explicit Motion Modeling for High-Quality Street Gaussian Splatting

### v1

@@ 4.1 Problem Formulation
Motion-aware Feature Encoding
@@
where $P$ is the number of frequency bands, enabling the network to capture both fine geometric details and global structures. For temporal information, we design an adaptive temporal embedding function:
@@
where $\mathbf{W}\in\mathbb{R}^{N_{max}\times D}$ is a learnable embedding matrix that captures motion patterns, and $N(i)$ progressively increases from $N_{min}$ to $N_{max}$ temporal samples during training iteration $i$, allowing the model to gradually capture finer temporal dynamics. For Gaussian-specific features, we assign a learnable latent embedding $\mathbf{z}_{k}\in\mathbb{R}^{D}$ to each Gaussian $k$, where $\mathbf{F}_{gauss}=\mathbf{z}_{k}$, enabling the model to learn and represent individual motion characteristics.
Table 1: Comparative performance of our framework and baseline approaches on the Waymo-NOTR dataset. The best performances are highlighted in bold, and the second-best are indicated with underlining. ↑ indicates higher is better, while ↓ indicates lower is better.
Scene Reconstruction Novel View Synthesis
Dataset Methods Full Image Vehicle Full Image Vehicle
PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑
D32 EmerNeRF 28.16 0.806 0.228 24.32 0.682 25.14 0.747 0.313 23.49 0.660
3DGS 28.47 0.876 0.136 23.26 0.716 25.14 0.813 0.165 20.48 0.753
MARS 28.24 0.866 0.252 23.37 0.701 26.61 0.796 0.305 22.21 0.697
S3Gaussian 30.69 0.900 0.121 26.23 0.804 26.62 0.824 0.159 22.61 0.681
Dual-scale Deformation Framework
@@
The final deformed parameters combine both coarse and fine scale predictions:
@@
Figure 3: Qualitative comparison on the self-supervised setting between S3Gaussian and S3Gaussian+ours EMD. We also visualize the error maps between the rendered images and ground truth to provide further insights.
@@ 4.2 Integration with Existing Frameworks
As mentioned in the introduction, current approaches for street scene reconstruction generally fall into supervised and self-supervised paradigms. Having formalized our EMD framework, we now demonstrate its integration into representative methods from both paradigms: StreetGaussian for supervised learning and S3Gaussian for self-supervised learning.
Self-supervised Integration: S3Gaussian
For self-supervised scenarios, we enhance S3Gaussian’s architecture with our motion-aware features. While S3Gaussian originally employs a Multi-head Gaussian Decoder for deformation prediction, we augment each Gaussian with our learnable embedding $\mathbf{z}_{k}$ and restructure its decoder into our dual-scale framework. Specifically, both coarse and fine stages predict deformations in position ($\Delta\mu$), opacity ($\Delta\alpha$), and spherical harmonics coefficients ($\Delta\mathbf{c}$), enabling more precise motion modeling through hierarchical refinement.
Supervised Integration: StreetGaussian
For supervised settings, StreetGaussian provides a framework that represents dynamic objects through tracked vehicle poses and object-specific Gaussians. Each object is characterized by tracked poses $\{R_{t},T_{t}\}_{t=1}^{N_{t}}$ that transform object Gaussians from local coordinates ($\mu_{o},R_{o}$) to world coordinates ($\mu_{w},R_{w}$). To enhance its motion modeling capability, we also augment each Gaussian with our learnable embedding $\mathbf{z}_{k}$ and apply temporal embedding only within each object’s tracked duration, effectively capturing object-specific temporal dynamics.
To address the challenge of noisy tracked poses while maintaining their geometric meaning, we incorporate our dual-scale framework into their pose optimization:
$R^{\prime}_{t}$ $=R_{t}(\Delta R_{t}^{c}+\Delta R_{t}^{f})$ (14)
@@
where $\Delta R_{t}^{c},\Delta T_{t}^{c}$ handle large pose corrections and $\Delta R_{t}^{f},\Delta T_{t}^{f}$ capture subtle adjustments. This decomposition allows the model to effectively correct tracking errors while preserving the physical meaning of the tracked poses. Similarly, we apply the dual-scale framework to appearance modeling, where spherical harmonics coefficients are refined through both coarse and fine stages, enabling more accurate dynamic appearance representation.
We adopt the same loss function as S3Gaussian and StreetGaussian to train the entire pipeline. For further training details, please refer to the supplementary materials.
@@ 5 Experiments
In this section, we first describe the datasets and evaluation metrics used in our experiments (Sec. 5.1). Then, we demonstrate the experimental results of our approach under both self-supervised and supervised settings (Sec. 5.2). Finally, we present comprehensive ablation studies to validate the effectiveness of individual components in our framework (Sec. 5.3).
@@ 5.1 Datasets and Metrics
Dataset.
We conduct extensive experiments on the Waymo Open Dataset (Sun et al. 2020) to comprehensively evaluate our method. To validate the effectiveness of our plug-and-play module, we select representative state-of-the-art methods from both supervised and self-supervised paradigms: StreetGaussian (Yan et al. 2024) and S3Gaussian (Huang et al. 2024). For self-supervised evaluation, we use the dynamic32 (D32) split introduced by EmerNeRF (Yang et al. 2023a), containing 32 sequences with vehicle motion, where each sequence consists of approximately 50-100 frames captured by three cameras under various conditions. For supervised evaluation, we follow the data split protocol established in StreetGaussian (Yan et al. 2024) to enable direct comparison.
Evaluation Metrics.
We employ comprehensive metrics to evaluate both reconstruction quality and novel view synthesis capability. For scene reconstruction, we use PSNR and SSIM to evaluate both full-scene and vehicle-specific reconstruction quality. Additionally, we compute LPIPS for perceptual quality assessment. Following previous protocols, we evaluate novel view synthesis on every 10th frame for self-supervised settings and every 4th frame for supervised settings.
@@ 5.2 Main Results
Self-supervised Performance
Tab. 1 presents the comparative results on the D32 split, where no 3D bounding box annotations were used. Our method significantly outperforms previous self-supervised approaches, achieving notable improvements in both full-scene and object-specific metrics. Specifically, we observe a substantial increase in PSNR for the full scene (32.50 vs. 30.69) and for vehicle-specific metrics (PSNR: 29.04 vs. 26.23) when compared to S3Gaussian. These results demonstrate the enhanced ability of our method to accurately model complex street scenes without explicit 3D box annotations. For novel view synthesis, our method continues to perform competitively, achieving a PSNR of 26.55, which reflects its robust generalization to previously unseen viewpoints.
In addition, we present a side-by-side visualization comparison between S3Gaussian and S3Gaussian+EMD in Fig. 3. The error maps, which compare the ground truth to the rendered results, clearly demonstrate that S3Gaussian+EMD outperforms S3Gaussian in modeling dynamic objects with varying motion speeds. S3Gaussian implicitly models dynamic vehicles, but it fails to capture changes in speed during motion and the differences in movement between vehicles, leading to blurred reconstructions. on the contrary, our method captures the distinct motion characteristics of different dynamic objects, leading to more accurate and consistent scene reconstructions. This emphasizes the effectiveness of Explicit Motion Decomposition (EMD) in modeling the motion of dynamic objects, improving the overall decomposition and photorealistic rendering of street scenes.
@@
Supervised Performance
Table 2: Comparative performance of our framework and baseline approaches on the StreetGaussian dataset. The best performances are highlighted in bold. ↑ indicates higher is better, while ↓ indicates lower is better.
@@
3DGS 29.64 0.918 0.117 21.25
NSG 28.31 0.862 0.346 24.32
MARS 29.75 0.886 0.264 26.54
EmerNeRF 30.87 0.905 0.133 21.67
StreetGaussian 34.61 0.938 0.079 30.23
StreetGaussian + Ours 36.03 0.949 0.067 29.81
To demonstrate the versatility of our framework, we also evaluate it in the supervised setting using the same scenes as StreetGaussian (Yan et al. 2024) (Tab. 2). When incorporating 3D bounding box supervision through our adaptive training scheme, our method achieves superior performance in full scene reconstruction, improving PSNR by 1.42dB (36.03 vs. 34.61), SSIM by 1.1% (0.949 vs. 0.938), and LPIPS by 15.2% (0.067 vs. 0.079). These results demonstrate that our approach effectively enhances the overall scene reconstruction quality while maintaining competitive performance in dynamic object modeling. Please refer to the supplementary for visualization on the supervised setting.
5.3 Ablation Studies
Table 3: Ablation study on the D32 dataset showing the impact of different components in our framework. All variants are evaluated using the self-supervised setting. The best performances are highlighted in bold. ↑ indicates higher is better, while ↓ indicates lower is better.
@@ 5.3 Ablation Studies
Both the coarse and fine deformation components are integral to the final performance, with their removal leading to considerable performance degradation. Specifically, excluding the coarse deformation component causes a significant 3.10 PSNR drop (29.40 vs. 32.50), suggesting that the coarse adjustments are vital for maintaining the overall scene structure. On the other hand, removing the fine deformation component results in a 0.036 LPIPS increase (0.118 vs. 0.082), implying that fine deformation refines the details and its absence worsens the perceptual quality. The ablation study demonstrates that our proposed motion-aware feature encoding and dual-scale deformation effectively model dynamic objects with varying motion speeds, which is crucial for enhancing the reconstruction quality of existing street Gaussians.
6 Limitation

### v2

@@ 4.1 Problem Formulation
4.1.1 Motion-aware Feature Encoding
@@
where $P=10$ is the number of frequency bands, enabling the network to capture both fine geometric details and global structures. For temporal information, we design an adaptive temporal embedding function:
@@
where $\mathbf{W}\in\mathbb{R}^{N_{max}\times D}$ is a learnable embedding matrix that captures motion patterns, and $N(i)$ progressively increases from $N_{min}=30$ to $N_{max}=150$ temporal samples during training iteration $i$, allowing the model to gradually capture finer temporal dynamics. $D=4$ is the temporal embedding dimension. During temporal embedding $\mathbf{W}$ optimization, we first downsample $\mathbf{W}$ through bilinear interpolation to create an intermediate embedding matrix of size $N(i)\times D$ at training iteration $i$, which can be formulated as:
$N(i)=N_{min}+(N_{max}-N_{min})\cdot\min(i,T)/T,$ (10)
where hyperparameter $T=25000$ controls the duration of progressive sampling refinement. Then, for each time stamp t, we perform grid sampling on this intermediate matrix with bilinear interpolation mode to obtain the corresponding temporal embedding vector $\mathbf{F}_{temp}(t)$. For Gaussian-specific features, we assign a learnable latent embedding $\mathbf{z}_{k}\in\mathbb{R}^{M}$ to each Gaussian $k$, where $\mathbf{F}_{gauss}=\mathbf{z}_{k}$ and $M=32$, enabling the model to represent individual motion characteristics.
4.1.2 Dual-scale Deformation Framework
@@
The final deformed parameters combine both coarse and fine-scale predictions:
@@ 4.2 Integration with Existing Frameworks
Current approaches for street scene reconstruction generally fall into supervised and self-supervised paradigms. Having formalized our EMD framework, we now illustrate its integration into representative supervised methods. Then we further extend EMD into supervised approaches.
4.2.1 Self-supervised Integration
For self-supervised street Gaussian splatting methods, we enhance S3Gaussian [16] and DeSiRe-GS [38] with our motion-aware techniques. S3Gaussian originally employs a HexPlane to model the dynamic and static element decomposition and utilizes a Multi-head Gaussian Decoder for deformation prediction. We augment each Gaussian with our learnable embedding $\mathbf{z}_{k}$ and restructure its decoder into our dual-scale framework. Specifically, both coarse and fine stages predict deformations in position ($\Delta\mu$), opacity ($\Delta\alpha$), and spherical harmonics coefficients ($\Delta\mathbf{c}$), enabling more precise motion modeling through hierarchical refinement. As for DeSiRe-GS built upon PVG [5], we also assign the learnable Gaussian embedding $\mathbf{z}_{k}$ into the framework. Then we apply the deformation for the position ($\Delta\mu$) and spherical harmonics coefficients ($\Delta\mathbf{c}$) with the proposed dual-scale deformation framework. We retain the same self-supervised deformation settings as those used in the baseline methods.
Table 1: S3Gaussian comparison: results on the Waymo Open dataset for scene reconstruction and novel view synthesis.
Scene Reconstruction Novel View Synthesis
Dataset Methods Full Image Vehicle Full Image Vehicle
PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑
Waymo-D32 EmerNeRF [61] 28.16 0.806 0.228 24.32 0.682 25.14 0.747 0.313 23.49 0.660
3DGS [19] 28.47 0.876 0.136 23.26 0.716 25.14 0.813 0.165 20.48 0.753
MARS [55] 28.24 0.866 0.252 23.37 0.701 26.61 0.796 0.305 22.21 0.697
S3Gaussian [16] 30.69 0.900 0.121 26.23 0.804 26.62 0.824 0.159 22.61 0.681
4.2.2 Supervised Extension
For supervised settings, though existing methods refine tracked 3D bounding boxes to eliminate rendering errors, we extend EMD into the refinement process, illustrating the need for explicit motion modeling. We select StreetGaussian [60] and OmniRe [6] as our baselines. StreetGaussian provides a framework that represents dynamic objects through tracked vehicle poses and object-specific Gaussians. Each object is characterized by tracked poses $\{R_{t},T_{t}\}_{t=1}^{N_{t}}$ that transform object Gaussians from local coordinates ($\mu_{o},R_{o}$) to world coordinates ($\mu_{w},R_{w}$). To enhance its motion modeling capability, we also augment each Gaussian with our learnable embedding $\mathbf{z}_{k}$ and apply temporal embedding only within each object’s tracked duration, effectively capturing object-specific temporal dynamics. Finally, we incorporate our dual-scale framework into the pose optimization:
$R^{\prime}_{t}$ $=\Delta R_{t}^{f}\cdot\Delta R_{t}^{c}\cdot R_{t}$ (13)
@@ 4.2.2 Supervised Extension
where $\Delta R_{t}^{c},\Delta T_{t}^{c}$ handle large pose corrections and $\Delta R_{t}^{f},\Delta T_{t}^{f}$ capture subtle adjustments. Similarly, we apply the dual-scale framework to appearance modeling, where spherical harmonics coefficients are refined through both coarse and fine stages.
For OmniRe, it further proposes non-rigid SMPL [27] nodes for human modeling. For the rigid nodes, we apply the same tracked box refinement in StreetGaussian into OmniRe. We further implement the dual-scale refinement to the SMPL model including pose parameters $\theta_{t}\in\mathbb{R}^{24\times 3\times 3}$ and shape parameters $\beta_{t}\in\mathbb{R}^{10}$:
$\theta^{\prime}_{t}$ $=\Delta\theta_{t}^{f}\cdot\Delta\theta_{t}^{c}\cdot\theta_{t}$ (14)
$\beta^{\prime}_{t}$ $=\beta_{t}+(\Delta\beta_{t}^{c}+\Delta\beta_{t}^{f}),$
For training, in addition to using the same loss function as the baselines, we implement a local smoothness regularization for the learnable Gaussian embeddings. Inspired by  [29], this regularization encourages neighboring Gaussians $i$ and $j$ to have similar representations:
$\mathcal{L}_{\mathbf{z}_{k}}=\frac{1}{d|\mathcal{U}|}\sum_{i\in\mathcal{U}}\sum_{j\in\text{KNN}_{i;d}}(e^{-\lambda_{w}\|\mu_{j}-\mu_{i}\|_{2}}\|\mathbf{z}_{k_{i}}-\mathbf{z}_{k_{j}}\|_{2}),$ (15)
where hyperparameters $\lambda_{w}=2000$ and $d=20$. KNN means the k-nearest-neighbors algorithm. We also regularize the predicted coarse and fine deformations, constraining their values to remain close to zero. For further training details, please refer to the supplementary materials.
Table 2: DeSiRe-GS comparison: results on the Waymo Open dataset and KITTI dataset for scene reconstruction and novel view synthesis.
Method Waymo Open KITTI
Scene Reconstruction Novel View Synthesis FPS Scene Reconstruction Novel View Synthesis FPS
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
S-NeRF [57] 19.67 0.528 0.387 19.22 0.515 0.400 0.0014 19.23 0.664 0.193 18.71 0.606 0.352 0.0075
StreetSurf [13] 26.70 0.846 0.3717 23.78 0.822 0.401 0.097 24.14 0.819 0.257 22.48 0.763 0.304 0.37
3DGS [19] 27.99 0.866 0.293 25.08 0.822 0.319 63 21.02 0.811 0.202 19.54 0.776 0.224 125
NSG [34] 24.08 0.656 0.441 21.01 0.571 0.487 0.032 19.19 0.683 0.189 17.78 0.645 0.312 0.19
Mars [55] 21.81 0.681 0.430 20.69 0.636 0.453 0.030 27.96 0.900 0.185 24.31 0.845 0.160 0.31
SUDS [50] 28.83 0.805 0.317 25.36 0.783 0.384 0.008 28.83 0.917 0.147 26.07 0.797 0.131 0.29
EmerNeRF [61] 28.11 0.786 0.373 25.92 0.763 0.384 0.053 26.95 0.828 0.218 25.24 0.801 0.237 0.28
PVG [5] 32.46 0.910 0.229 28.11 0.849 0.279 50 32.83 0.937 0.070 27.43 0.896 0.114 59
DeSiRe-GS [38] 33.61 0.919 0.204 29.75 0.878 0.213 36 33.94 0.949 0.040 28.87 0.901 0.106 41
DeSiRe-GS + ours 34.15 0.948 0.130 29.91 0.880 0.190 32 34.13 0.954 0.040 29.05 0.904 0.094 32
Figure 3: Visualization comparison on the self-supervised setting between S3Gaussian and S3Gaussian+ours EMD. We also visualize the error maps between the rendered images and ground truth to provide further insights. Please refer to appendix for more visualization.
@@ 5.1 Datasets and Metrics
Dataset and Baselines. We conduct extensive experiments on the Waymo Open Dataset [44] and KITTI Dataset [12] to comprehensively evaluate our method. For self-supervised setting, we select representative state-of-the-art methods including S3Gaussian [16] and DeSiRe-GS [38]. To compare with S3Gaussian, we use the dynamic32 (D32) split with 3 frontal cameras of Waymo dataset introduced by EmerNeRF [61] and benchmark S3Gaussian + EMD with vanilla 3DGS [19] and MARS [55] on scene reconstruction and novel view synthesis. To compare with DeSiRe-GS which follows PVG [5], we use the same subset in PVG, including 4 Waymo scenes and 3 KITTI scenes. We implement baselines including S-NeRF [57], StreetSurf [13], NSG [35] and SUDS [50] on scene reconstruction and novel view synthesis. For the supervised setting, we select state-of-the-art approaches including StreetGaussian [60] and OmniRe [6]. We select the same Waymo subset from StreetGaussian and OmniRe and conduct experiments on novel view synthesis. It should be noted that StreetGaussian only conducts experiments with one frontal camera in its published version. Thus, we further implement experiments on three frontal cameras for a side-by-side comparison.
Evaluation Metrics. Following previous protocols, we use PSNR and SSIM to evaluate the pixel-level reconstruction quality. Additionally, we compute LPIPS for perceptual quality assessment. We also report FPS to access inference speed in the comparison between DeSiRe-GS, which refers to frames per second. In addition, we employ the FID metric [15, 65] for novel trajectory synthesis evaluation, which quantifies differences in feature distribution between rendered novel trajectory images and original trajectory images.
@@ 5.2 Main Results
5.2.1 Self-supervised Performance
Tab. 1 presents the comparative results on the Waymo-D32 split, where no 3D bounding box annotations were used. Our method significantly outperforms previous self-supervised approaches, achieving notable improvements in both full-scene and object-specific metrics. These results demonstrate the enhanced ability of our method to accurately model complex motions in street scenes without explicit 3D box annotations. S3Gaussian models the entire scene holistically but lacks the ability to capture the diverse motion patterns in dynamic street scenes, whereas EMD effectively compensates for this shortcoming. For novel view synthesis, our method continues to perform competitively, which reflects its robust generalization to previously unseen viewpoints.
Tab. 2 compares DeSiRe-GS with our proposed EMD. The results show that EMD achieves state-of-the-art performance across various experimental settings. DeSiRe-GS relies on dynamic masks to identify foreground objects, thereby overlooking motion patterns in street scenes. EMD boosts the rendering quality of DeSiRe-GS with only a slight reduction in inference speed. These quantitative results effectively demonstrate that EMD is well-suited for current self-supervised street Gaussian methods.
In addition, we present a side-by-side visualization comparison between S3Gaussian and S3Gaussian+EMD in Fig. 3. The error maps, which compare the ground truth to the rendered results, clearly demonstrate that S3Gaussian+EMD outperforms S3Gaussian in modeling dynamic objects with varying motion speeds. S3Gaussian implicitly models dynamic vehicles, but it fails to capture changes in speed during motion and the differences in movement between vehicles, leading to blurred reconstructions. on the contrary, our method captures the distinct motion characteristics of different dynamic objects, leading to more accurate and consistent scene reconstructions. This emphasizes the effectiveness of Explicit Motion Decomposition (EMD) in modeling the motion of dynamic objects, improving the overall decomposition and photorealistic rendering of street scenes. For more visualization videos, please watch the webpage Sec.B in the supplementary materials.
@@ 5.2.1 Self-supervised Performance
Table 3: StreetGaussin comparison: novel view synthesis results on the Waymo Open dataset with one and three frontal cameras.
@@
One camera setting
3DGS [19] 29.64 0.918 0.117 21.25
NSG [34] 28.31 0.862 0.346 24.32
MARS [55] 29.75 0.886 0.264 26.54
EmerNeRF [61] 30.87 0.905 0.133 21.67
StreetGaussian [60] 34.61 0.938 0.079 30.23
StreetGaussian + Ours 35.41 0.942 0.070 30.96
Three camera setting
StreetGaussian 29.70 0.858 0.149 26.72
StreetGaussian + Ours 29.84 0.869 0.145 26.83
Table 4: OmniRe comparison: novel view synthesis results on the Waymo Open dataset with three frontal cameras.
Methods Full Image Human Vehicle
PSNR↑ SSIM↑ PSNR↑ SSIM↑ PSNR↑ SSIM↑
EmerNeRF [61] 29.67 0.883 20.32 0.454 22.07 0.609
3DGS [19] 25.57 0.906 16.62 0.387 16.00 0.407
DeformGS [63] 27.72 0.922 17.30 0.426 18.91 0.530
PVG [5] 30.19 0.919 21.30 0.567 22.28 0.679
HUGS [67] 27.65 0.914 15.99 0.378 23.27 0.748
StreetGS [60] 28.54 0.928 16.55 0.393 26.71 0.846
OmniRe [6] 32.57 0.942 24.36 0.727 27.57 0.858
OmniRe + Ours 33.89 0.958 25.97 0.742 27.82 0.859
5.2.2 Supervised Performance
To demonstrate the flexibility of EMD, we further extend it to enhance motion modeling in supervised methods. Table 3 and Table 4 show novel view synthesis comparisons between StreetGaussian and OmniRe, respectively. While these supervised methods employ 3D box refinement to mitigate tracking errors, the results indicate that EMD serves as a valuable complement to this process. In the OmniRe comparison, EMD also benefits non-rigid objects such as humans, whose motion patterns differ significantly from those of vehicles. Due to the space limitation, please refer to the supplementary for visualization on the supervised setting.
Table 5: Novel trajectory synthesis on Waymo dataset.
Method FID↓
0.5m 1.0m 1.5m
S3Gaussian 83.48 110.11 134.38
S3Gaussian + ours 45.11 70.26 90.20
Figure 5: Visualization for novel trajectory synthesis (0.5m offset). Please watch the webpage in the supplementary materials for more results.
5.3 Novel Trajectory Synthesis
Previous novel view synthesis evaluations are limited to interpolated views along the original camera trajectory, which fail to assess the exact simulation performance of the model. Therefore, we compare the FID performance of self-supervised methods and our approach under novel trajectory synthesis. Specifically, we shift the original camera trajectory to the left and right by different offsets (0.5 m, 1.0 m, 1.5 m) and render images along these new trajectories. The results are presented in Tab. 5 and one visualization sample is shown in Fig. 5. These findings demonstrate that EMD enhances the accurate modeling of dynamic objects by learning diverse motion patterns, which benefits lane change scenarios. For novel trajectory synthesis videos, please refer to the supplementary materials.
Table 6: Ablation study on the 32 dynamic scenes Waymo subset showing the impact of different components in our framework. All variants are evaluated in self-supervised scene reconstruction.
@@ 5.4 Ablation Studies
Both the coarse and fine deformation components are integral to the final performance, with their removal leading to performance degradation. Specifically, excluding the coarse deformation component causes a significant 3.10 PSNR drop (29.40 vs. 32.50), suggesting that the coarse adjustments are vital for maintaining the overall scene structure. On the other hand, removing the fine deformation component results in a 0.036 LPIPS increase (0.118 vs. 0.082), implying that fine deformation refines the details and its absence worsens the perceptual quality. The ablation study demonstrates that our proposed motion-aware feature encoding and dual-scale deformation effectively model dynamic objects with varying motion speeds, which is crucial for enhancing the reconstruction quality of existing street Gaussians.
6 Conclusion
In this paper, we present EMD, the first plug-and-play module that effectively handles varying motion speeds in street scene reconstruction. By introducing motion-aware feature encoding and dual-scale deformation modeling, our approach successfully captures complex motion patterns in real-world scenarios. Comprehensive experiments on the Waymo and KITTI datasets demonstrate that EMD achieves the state-of-the-art novel view synthesis quality for self-supervised frameworks. EMD can be extended into supervised street Gaussian splatting methods, which opens up new possibilities for high-quality driving scene reconstruction.
Acknowledgments.
This work was supported by the National Science and Technology Major Project (No. 2022ZD0117800).
Supplementary Material
Appendix A Overview
The supplementary material includes the subsequent components.
  • Additional Visualization Videos
  • Implementation Details
    • Training Schemes
    • Training Details
    • Parameters and Efficiency
  • Parameter Sensitivity
Appendix B Additional Visualization Videos
Please double-click the “Demo Webpage-Please wait until loaded.html" file and open it in your browser. This offline webpage contains videos covering the following experiments:
  • Self-supervided Comparison
  • Box Supervised Comparison
  • Novel Trajectory Synthesis
  • Temporal Embedding Evolution
Due to the numerous videos, please wait for the webpage until loaded.
Appendix C Implementation Details
C.1 Training Schemes
LiDAR Prior Initialization. To initialize the positions of the 3D Gaussians, we leverage the LiDAR point cloud captured by the vehicle instead of using the original SFM [42] point cloud to provide a better geometric structure. To reduce model size, we also downsample the entire point cloud by voxelizing it and filtering out points outside the image. For colors, we initialize them randomly.
Optimization Objective. Following Street Gaussian, we introduce the sky supervision loss $L_{sky}$ into the original loss function proposed by S3Gaussian. Subsequently, we get a composed training loss function which can impose various constraints to our model.
$\begin{split}\mathcal{L}=\mathcal{L}_{color}&+\lambda_{depth}\mathcal{L}_{depth}+\lambda_{feat}\mathcal{L}_{feat}\\ &+\lambda_{tv}\mathcal{L}_{tv}+\lambda_{sky}\mathcal{L}_{sky}+\lambda_{reg}\mathcal{L}_{reg}\end{split}$ (16)
Here, $\mathcal{L}_{depth}$ is the mean square error (MSE) loss between the rendered depth map and the estimated depth map from the LiDAR point cloud, which aids in supervising the expected position of 3D Gaussians. $\mathcal{L}_{feat}$ is also the L2 loss of semantic features to reduce the gap between both planes. $\mathcal{L}_{tv}$ is a total-variational loss based on grids to make rendered objects smoother. $\mathcal{L}_{color}$ is the main loss to give constraints to the reconstruction process formulated by:
$\mathcal{L}_{color}=\mathcal{L}_{rgb}+\lambda_{ssim}\mathcal{L}_{sim}$ (17)
Furthermore, $\mathcal{L}_{reg}$ is organized as:
$\mathcal{L}_{reg}=\mathcal{L}_{z_{k}}+\mathcal{L}_{\Delta}$ (18)
where $\mathcal{L}_{z_{k}}$ is local smoothness regularization for Gaussian embeddings in the method section. $\mathcal{L}_{\Delta}$ represents a combination of regularization for coarse and fine deformations, restricting their values near zero. We also detail the coefficients for loss in Tab. 7.
Table 7: Loss function coefficients
$\lambda_{depth}$ $\lambda_{feat}$ $\lambda_{feat}$ $\lambda_{tv}$ $\lambda_{sky}$ $\lambda_{reg}$
0.5 0.1 0.1 0.1 0.1 0.01
Table 8: Parameter sensitivity analysis on the D32 dataset, highlighting the effect of varying the dimensions of Gaussian embeddings $\mathbf{z}_{k}$ and temporal embeddings $\mathbf{z}_{w}$. All experiments are conducted in the self-supervised setting. Best performances are highlighted in bold. ↑ indicates higher is better, while ↓ indicates lower is better. We also include the changes in model parameters relative to the adopted setting.
$\mathbf{z}_{k}$/$\mathbf{z}_{w}$ Parameters Full Image Vehicle
PSNR↑ SSIM↑ LPIPS↓ PSNR↑
32/4 / 32.50 0.933 0.082 29.04
128/4 +14400 32.22 0.925 0.086 29.05
8/4 -3600 31.25 0.910 0.128 27.75
32/4 / 32.50 0.933 0.082 29.04
32/16 +14.42M 32.38 0.930 0.081 29.01
32/1 -3.60M 30.55 0.898 0.136 27.04
C.2 Training Details
For S3Gaussian, we train the entire pipeline for 50,000 iterations using the Adam optimizer. Following the original S3Gaussian setup, we perform a warm-up phase for each scene, employing 5,000 iterations to train a coarse representation using vanilla 3D Gaussians. After this warm-up phase, we integrate the proposed dual-scale deformation network, which is jointly optimized with the HexPlane. To implement a coarse-to-fine training strategy, temporal embeddings $N(i)$ are progressively increased from $N_{min}$ to $N_{max}$ in 20,000 iterations, allowing for the gradual motion modeling of objects. Since S3Gaussian is evaluated on 50 frames per clip for each scene, we ensure a fair comparison by conducting all self-supervised validation experiments on the first clip of 32 dynamic scenes. Other configurations, including the detailed setup of the HexPlane and learning rates, are kept consistent with the S3Gaussian. For StreetGaussian, the entire method is trained for 30,000 iterations on a subset of eight selected scenes from the StreetGaussian dataset. Unlike the self-supervised method, we bind the proposed EMD to the vehicle Gaussians in each scene. Temporal embeddings are applied based on the time each vehicle appears within the scene. All other settings, including detailed configurations, remain consistent with those described in StreetGaussian. All experiments are conducted on a single NVIDIA A800 GPU.
Appendix D Parameter Sensitivity
To analyze the sensitivity of model performance to the dimensions of Gaussian embeddings $\mathbf{z}_{k}$ and temporal embeddings $\mathbf{z}_{w}$ (derived from the learnable embedding matrix $\mathbf{W}$), we conduct experiments by varying these dimensions. In the original setup, $\mathbf{z}_{k}$ is set to 32 and $\mathbf{z}_{w}$ to 4. Tab. 8 summarizes the results, demonstrating how these changes influence performance under the self-supervised setting.
The results reveal that reducing the embedding dimensions leads to significant performance degradation. This is primarily due to the reduced capacity to effectively model the motion of dynamic objects, which is crucial for high-quality reconstruction. On the other hand, increasing the embedding dimensions offers only marginal performance improvements. However, due to the large number of Gaussians in the driving scenes, the higher embedding dimensions result in a substantial increase in the total number of model parameters, leading to a higher computational cost. These findings highlight the trade-off between embedding dimension size and overall model efficiency. While lower dimensions compromise the ability to capture dynamic motion, higher dimensions introduce considerable overhead without proportional gains in performance. The adopted embedding configuration achieves a good balance, maintaining strong performance while keeping the parameter count manageable.
Appendix E Limitation